For a Bose-Einstein system, the expression for the thermodynamic probability is

W=(3.14)((gi+Ni-1)!/(Ni!(gi-1)!))

Derive an expression for the Bose-Einstein distribution function.


1
Expert's answer
2021-03-14T19:15:23-0400
"f=(e^{\frac {E-\mu _\epsilon}{k \tau}}-1)^{-1}"

"\mu_\epsilon=\frac {dU}{dn}=0"

"f=(e^{\frac {h\nu}{kT}} -1)^{-1}"